# The Dark Side of NFTs: A Large-Scale Empirical Study of Wash Trading

### v1

@@ VIII-B Implications for NFT Marketplace
IX RELATED WORK
IX-A Data collection for NFT research.
NFT researchers mainly collect data in an external way. Von et al. [13] utilized OpenSea API to collect NFTs’ events and adopted Coingecko API to retrieve the historic USD prices for crypto. White et al. [46] utilized a moving window of Unix timestamps to retrieve 5,252,252 sale events via OpenSea API. Apart from the API access, Das et al. [14] collected more data through web scraping. However, few researchers delve deeper and discuss the completeness and consistency of the data provided by the marketplace, e.g., OpenSea. As we demonstrate in section IV, Address inconsistency, Timestamp error and Fake events all occur in the UI and database of OpenSea. To the best of our knowledge, we are the first paper to raise and address the inconsistency problem for NFTs’ event sequences, providing a robust foundation for further research.
IX-B Detection for NFT wash trading.
Several studies have been published to detect wash trading behavior for crypto assets. In terms of Bitcoin, Aloosh et al. [47] provided direct evidence for ‘fake volume’ that occurred in cryptocurrency exchanges through trading records leaked by hackers. For NFTs, current researchers pay more attention to NFT TXN networks rather than discussing different types of wash trading, including evidence from block/ERC-20 TXNs and Private Trading. For example, Von et al. [13] presented methods to identify the graph patterns for wash trading, indicating that 2.04% of sale events are suspicious. Based on that, La et al. [27] explored more suspicious graph patterns. Das et al. [14] focused on SCC and WCC to discover the wash trading behavior. Wen et al. [17] provided a novel visualization method for identification. In addition to La et al. [27] who analyzed the profitability of wash trading, we are the first to place a broader discussion on forms/impacts of NFT wash trading, instead.
X Conclusions
This paper introduces currently the most comprehensive understanding of NFT wash trading. We first organize a robust dataset containing 8,717,031 transfer events and 3,830,141 sale events from 2,701,883 NFTs of the 285 most popular collections. Then, we propose heuristic algorithms to identify NFT wash trading/traders. At last, we conduct extensive experiments to measure the financial impact/trend/market liquidity of NFT wash trading, and present six findings from the marketplace design, profitability, NFT project design, payment token, user behavior, and NFT ecosystem.

### v2

@@ 8.2. Implications for NFT Marketplace
9. RELATED WORK
9.1. Data collection for NFT research.
NFT researchers mainly collect data in an external way. Von Wachter et al. (von Wachter et al., 2022) utilized OpenSea API to collect NFTs’ events and adopted Coingecko API to retrieve the historic USD prices for crypto. White et al. (White et al., 2022) utilized a moving window of Unix timestamps to retrieve 5,252,252 sale events via OpenSea API. Apart from the API access, Das et al. (Das et al., 2021) collected more data through web scraping.
9.2. Detection for NFT wash trading.
Several studies have been published to detect wash trading behavior for crypto assets. In terms of Bitcoin, Aloosh et al. (Aloosh and Li, 2019) provided direct evidence for ‘fake volume’ that occurred in cryptocurrency exchanges through trading records leaked by hackers. For NFTs, current researchers pay more attention to NFT TXN networks rather than discussing different types of wash trading, including evidence from block/ERC-20 TXNs and Private Trading. For example, von Wachter et al. (von Wachter et al., 2022) presented methods to identify the graph patterns for wash trading, indicating that 2.04% of sale events are suspicious. Based on that, La et al. (La Morgia et al., [n. d.]) explored more suspicious graph patterns. Das et al. (Das et al., 2021) focused on SCC and WCC to discover the wash trading behavior. Wen et al. (Wen et al., 2023) provided a novel visualization method for identification. In addition to La et al. (La Morgia et al., [n. d.]) who analyzed the profitability of wash trading, we are the first to place a broader discussion on forms/impacts of NFT wash trading, instead.
10. Conclusions
This paper introduces currently the most comprehensive understanding of NFT wash trading. We first organize a dataset containing 8,717,031 transfer events and 3,830,141 sale events from 2,701,883 NFTs of the 285 most popular collections. Then, we propose heuristic algorithms to identify NFT wash trading/traders. At last, we conduct extensive experiments to measure the financial impact/trend/market liquidity of NFT wash trading, and present six findings from the marketplace design, profitability, NFT project design, payment token, user behavior, and NFT ecosystem.
Acknowledgements.
This research/project is supported by the Open Research Fund of The State Key Laboratory of Blockchain and Data Security, Zhejiang University), the National Natural Science Foundation of China (62302534).